My daughter has this question in a maths book
A) Dropped from 1 meter onto a hard surface, a golf ball bounced 49cm; calculate e, the coefficient of restitution, or bounciness. Of the ball.
B) The ball weighed 46g, and was hit by a 200g club-head travelling at45m/s.
Calculate vm/s, the speed of the ball as it left the club-head, to the nearest m/s,
I have worked this out to be 200*45(1+1.43)
200+46
C) Find the length L yards of the drive in b, to the nearest yard.

The answer is suppose to be A) 0.7 B) 62 m/s C) 228 yards.
I cant see were I’m going wrong please help

Answer by stanbon(75887)

My daughter has this question in a maths book
A) Dropped from 1 meter onto a hard surface, a golf ball bounced 49cm; calculate e, the coefficient of restitution, or bounciness. Of the ball.
---
I was able to find an online site for coefficient of restitution.
Cr = sqrt[h/H] where h is bounce height and H is drop height
---
Cr = sqrt[49/100] = 7/10 = 0.7
--------
Note: I cannot find an online site that explains club head speed
and distance calculations.
Maybe your daughter's book has formulas you can use and/or
calculation examples.
----------
Cheers,
Stan H.
-----------
B) The ball weighed 46g, and was hit by a 200g club-head travelling at45m/s.
Calculate vm/s, the speed of the ball as it left the club-head, to the nearest m/s,
I have worked this out to be 200*45(1+1.43)
200+46
C) Find the length L yards of the drive in b, to the nearest yard.

The answer is suppose to be A) 0.7 B) 62 m/s C) 228 yards.
I cant see were I’m going wrong please help